# Understanding the Effectiveness of Large Language Models in Detecting Security Vulnerabilities

### v2

@@ 2.2. Metrics
F1 score: The F1 score is a harmonic mean of precision and recall. It is computed as:
2.3. Large Language Models
We choose the most popular state-of-the-art pre-trained Large Language Models (LLMs) for our evaluation. We choose OpenAI models: GPT-4 (gpt-4) and GPT-3.5 (gpt-3.5-turbo). GPT-3.5 allows up to 4096 input tokens while GPT-4 (which presumably is much larger) allows up to 8192 tokens in the input prompt. Since these models are closed-source, we also evaluate CodeLlama models (Rozière et al., 2023), which were recently open-sourced by Meta. We select three versions of these models: CodeLlama-7B (CL-7B), CodeLlama-13B (CL-13B), and CodeLlama-34B (CL-34B), containing 7 billion, 13 billion, and 34 billion parameters, respectively. We use the Hugging Face APIs (Hugging Face, 2023) to access CodeLlama models. We use the “Instruct” version of CodeLlama models—these models are fine-tuned to follow user instructions and hence can better adapt to specific reasoning tasks. Table 3 presents LLM details.
| Model Class | Model Version | Size | Context Window |
@@ 2.3. Large Language Models
| GPT-3.5 | gpt-3.5-turbo-0613 | N/A | 4k |
| CodeLlama-34B | CodeLlama-34B-Instruct | 34B | 16k |
| CodeLlama-13B | CodeLlama-13B-Instruct | 13B | 16k |
| CodeLlama-7B | CodeLlama-7B-Instruct | 7B | 16k |
2.4. Prompting Strategies for LLMs
We explore various prompting strategies that can assist LLMs in predicting if a given code snippet is vulnerable. The LLMs discussed in this study support chat interactions with two major types of prompts: the system prompt and the user prompt. The system prompt can be used to set the context for the entire conversation while user prompts can be used to provide specific details throughout the chat session. We include a system prompt at the start of each input to describe the task and expected structure of the response. Since persona assignment has been shown to improve the performance of GPT-4 on specialized tasks (Salewski et al., 2024), we add the line “You are a security researcher, expert in detecting security vulnerabilities” at the start of every system prompt to assign a persona of a Security Researcher to the model. The system prompt for all experiments ends with the statement “Provide response only in the following format:” followed by an expected structure of the response from the model. The system prompt is followed by a user prompt that varies across the various prompting strategies. In all our experiments, we incorporate the target code snippet into the user prompt without any changes.
We construct different prompting strategies to study the reasoning abilities of LLMs:
2.4.1. Basic prompt
We design a very simple prompt (shown in Listing 4 in the appendix) with the goal of understanding if the model can take a target code snippet as input and detect if it is vulnerable and determine the correct CWE as well. The prompt begins with the message “Is the following code snippet prone to any security vulnerability?” followed by the code snippet.
| OWASP | Juliet | Juliet | CVEFixes | CVEFixes | Total | |
@@ 2.4.2. CWE specific prompt
The CWE specific prompt is presented in Listing 5 in the appendix. This prompt is similar to the Basic prompt except that it asks the model to predict if the given code snippet is vulnerable to a specific target CWE. Hence, the user prompt starts with “Is the following code snippet prone to ¡CWE¿?” followed by the code snippet. The <CWE> placeholder here contains both the ID of the CWE and the name. For instance, for CWE-22, the user prompt would start with “Is the following code snippet prone to CWE-22 (Improper Limitation of a Pathname to a Restricted Directory (Path Traversal))?” followed by the target code snippet.
2.4.3. Dataflow analysis-based prompt
@@
In addition to the straightforward one-step analysis in the previous two prompts, we also study whether providing specific step-by-step analysis instructions can help the LLMs do better reasoning over code and make better predictions. Dataflow analysis is used by several static analysis tools to infer if there exists an unsanitized path from a source to a target node. Further, prior literature has shown step-by-step instructions can often elicit better reasoning from LLMs (Wei et al., 2022b). Motivated by these observations, we designed the CWE-DF prompt (shown in Listing 6 in appendix) that prompts the model to simulate a source-sink-sanitizer-based dataflow analysis on the target code snippet before predicting if it is vulnerable. Naturally, compared to the other prompts, this prompt generates many more tokens and is hence more costly. We provide the full prompts in Appendix A.2.
2.4.4. Other prompting strategies
@@
We also tried other prompting strategies such as Few-shot prompting and Chain-of-thought prompting. In the few-shot prompting setup, we include two examples of the task (one with a vulnerability and one without) in the CWE specific prompt before providing the target code snippet. Few-shot prompting reported poorer results than the base CWE specific prompt while requiring more tokens. Our analysis of the few-shot prompts suggests that providing more examples might not be a useful strategy for vulnerability detection. It might be more useful to use prompts that instead elicit reasoning or explanations of some kind before detecting if the given snippet is vulnerable. With Chain-of-thought prompting, we explicitly ask the model to provide a reasoning chain before the final answer by adding a “Let’s think step-by-step” statement at the end of the CWE specific prompt. This setup did not yield better results than the Dataflow analysis-based prompt. Moreover, the reasoning chains obtained by Chain-of-thought prompting (both zero-shot and few-shot) were not as elaborate as those from the Dataflow analysis-based prompt thus limiting the ease of debugging. We provide the details of this experiment in Appendix A.3.
2.5. Dataset Processing and Selection
@@
We perform a data processing and cleaning step for each dataset before evaluating them with LLMs. We remove or anonymize information such as commits, benchmark IDs, or vulnerability names that may provide obvious hints about the vulnerability. We skip benchmarks that are spread across multiple files, due to limitations of prompt size. Table 4 presents the details of our selection stages.
Due to the prohibitively large cost of running pre-trained LLMs, we select a subset of samples from the original datasets. We select samples corresponding to vulnerability types (or CWEs) listed in MITRE’s Top 25 Most Dangerous Software Weaknesses (MITRE Top 25 CWEs, 2023). Due to prompt size limitations, we filter out code snippets larger than 2048 tokens. Finally, because the datasets are quite large, we further randomly select 500 vulnerable and 500 non-vulnerable samples per dataset. Finally, we end up with 5000 samples across five datasets. We provide more details for each dataset in Appendix A.1.
2.6. Experimental Setup
@@
Experiments with GPT-3.5 and GPT-4. We use the OpenAI public API to perform the experiments with GPT-3.5 and GPT-4. We use the ChatCompletions API endpoint to query the models with the prompts discussed in Section 2.4. We set the sampling temperature to 0 for obtaining deterministic predictions, the maximum number of tokens to 1024, and use the default values for all other parameters. In all our experiments, we use the top-1 prediction from the models.
Experiments with CodeLlama. We run all CodeLlama experiments on two sets of machines: one with Intel Xeon machine, with 40 CPUs, four GeForce RTX 2080 Ti GPUs, and 750GB RAM, and another Intel Xeon machine with 48 CPUs, 8 A100s, and 1.5T RAM. Similar to GPT-4, we set the temperature to 0, the maximum tokens to 1024, and use top-1 prediction for evaluation.
3. Results
3.1. RQ1: Effectiveness of LLMs
@@
We evaluate the performance of pre-trained LLMs on five open-source datasets discussed in Section 2.1. presents the best accuracy and F1 scores (across prompts) of GPT-4, GPT-3.5, and CodeLlama models (CL-7B, CL-13B, CL-34B) on all datasets. The more detailed metrics for all prompts are presented in Appendix A.4.
Modest Vulnerability Detection Performance Across LLMs. The average of the highest accuracies across all datasets is 60.6% with the maximum being 77% for CL-7B (with Basic prompt) on the Juliet Java dataset. In terms of F1 score, GPT-4 (with CWE-DF) performs the best on the synthetic datasets: OWASP (F1: 0.69), Juliet Java (F1: 0.76), and Juliet C/C++ (F1: 0.70). We observe that CodeLlama models, despite being much smaller than GPT-4, are competitive on most datasets, even with the Basic prompt. CL-34B reports the highest F1 scores (F1: 0.67) on the real-world datasets, CVEFixes Java and CVEFixes C/C++, with the CWE-DF prompt.
Performance does not improve with scale. While GPT-4 performs strictly better than GPT-3.5 on 3 of the 5 datasets, the CodeLlama models roughly exhibit similar performance. Surprisingly, CL-7B performs better than the larger LLMs on the Juliet Java and Juliet C/C++ datasets even with the Basic prompt. Listing LABEL:lst:juliet1724 presents a representative example where only CL-7B correctly predicts that an integer overflow vulnerability (CWE-190) cannot occur in the given context. Since the data variable can only contain value 2, adding 1 to it in Line LABEL:lst:line:intoverflow:add cannot cause an integer overflow. CL-7B recognizes this and predicts that the snippet is not vulnerable. CL-13B and GPT-4 (CWE-DF) ignore the hard-coded values and predict that the snippet is vulnerable, indicating that smaller LLMs may sometimes perform better than larger models.
@@
3.2. RQ2: Comparison of Prompting Strategies
@@
Figure 2 presents the accuracy and F1 scores (averaged across all LLMs) of the three prompting strategies across all datasets. Overall, the three prompts perform similarly in terms of accuracy. CWE-DF performs better than CWE and Basic prompt in terms of F1 score on the real-world datasets (CVEFixes C/C++ and CVEFixes Java).
We next highlight qualitative differences between various prompts:
The Basic prompt often detects incorrect vulnerability classes. Table 5 presents the percentage of samples predicted as vulnerable with the Basic prompt where the predicted CWE is correct, averaged over all datasets with the same programming language. We can observe that all models predict an incorrect vulnerability in ¿ 60% and ¿ 53% of all Java and C/C++ samples predicted as vulnerable respectively. This suggests that the Basic prompt detects incorrect CWEs in roughly half of the cases.
| Language (Avg.) | GPT-4 | GPT-3.5 | CL-34B | CL-13B | CL-7B |
| Java | 0.41 | 0.34 | 0.37 | 0.38 | 0.39 |
| C/C++ | 0.29 | 0.31 | 0.33 | 0.35 | 0.47 |
@@
Specifying the CWE in the prompt reduces false alarms. In 16 out of the 25 model-dataset combinations, the CWE specific prompt improves or retains both the accuracy and F1 score over the Basic prompt. GPT-4 with the CWE specific prompt on Juliet Java reports a 13% higher accuracy and a 0.1 higher F1 score than those with the Basic prompt. We manually inspect 10 vulnerable and 10 non-vulnerable samples from this dataset where the CWE specific prompt is correct and the Basic prompt is not. We find that not including the CWE in the Basic prompt results in predictions discussing other incorrect CWEs. The Basic prompt incorrectly predicts that the 10 non-vulnerable samples are vulnerable to other CWEs of which we only find 3 to be plausible but unlikely due to the context. For example, it predicts that a value read from an input stream can be vulnerable if not validated (CWE-20) but this value is not used in a vulnerable context. Of the 10 vulnerable samples, Basic prompt predicts a plausible CWE for one sample while incorrectly predicting that the remaining 9 samples are not vulnerable. On the other hand, the CWE specific prompt also provides useful high-level explanations for why the snippet is vulnerable/not vulnerable in 18 / 20 samples. The 2 incorrect explanations are artifacts of faulty reasoning or hallucination: for example, an Integer Overflow due to addition to INT_MAX in the function is incorrectly attributed to subtracting from INT_MIN in the explanation.
@@
Dataflow analysis identifies CWE-relevant textual cues and provides more elaborate explanations. The Dataflow analysis-based prompt (CWE-DF) prompt performs better than CWE specific prompt both in terms of accuracy and F1 score in 12 out of 25 model-dataset combinations and reports a higher F1 score in 23 out of 25 cases. The most significant improvement with the CWE-DF prompt over the CWE specific prompt is for the CVEFixes Java dataset and GPT-3.5 with 7% higher accuracy and 0.4 higher F1 score. We manually inspect 10 vulnerable and 10 non-vulnerable samples from this dataset where only the CWE-DF prompt is correct. We find that the CWE-DF prompt leverages textual cues for sanitization (e.g., csrftokenhandler() suggests protection from CSRF) while the CWE specific prompt does not. The CWE-DF prompt also assumes that any external parameter or call source is tainted by design and predicts that the sample is vulnerable more often (as is evident from higher F1). Further, the analysis from CWE-DF prompt is more detailed and useful in localization than the explanations from CWE specific prompt.
@@
3.3. RQ3: Performance of LLMs across CWEs
@@
We next evaluate how the LLMs perform on different classes of security vulnerabilities (CWEs). For each dataset and model, we consider the best-performing prompt for the analysis and only report CWEs with at least 10 samples. Because the CWE-wise distribution of vulnerable and non-vulnerable samples can be imbalanced, we compute balanced accuracy for each CWE (for ease of presentation, we refer to it as accuracy henceforth in this section). Figure LABEL:fig:java_cwe_distribution presents the CWE-wise distribution of accuracies on the OWASP, Juliet Java, and CVEFixes Java datasets. Figure 3 reports the accuracies on the Juliet C/C++ and CVEFixes C/C++ datasets respectively. For each model, we consider the prompt with the best accuracy on the dataset and break ties using the F1 score.
LLMs perform well on vulnerabilities that do not require additional context. We find that there are certain CWEs where multiple LLMs perform better on: Out-of-bounds Read / Write (CWE-125, CWE-787), Null pointer dereference (CWE-476) and Integer Overflow (CWE-190). The higher performance on these vulnerabilities can be attributed to the fact that these are fairly self-contained and little additional context is needed to detect them. Concretely, GPT-3.5 consistently performs well on NULL Pointer Dereference across all datasets with accuracies of 80.8%, 87.9% and 60.5% on the Juliet C/C++, Juliet Java and CVEFixes C/C++ datasets respectively. GPT-4 also performs consistently better on NULL Pointer Dereference and Integer Overflow on the synthetic datasets with accuracies of 62.1% and 68.4% on Juliet C/C++ and 78.5% and 73% on Juliet Java respectively. CL-7B reports high accuracies on Integer Overflow on the synthetic datasets with accuracies of 81% on Juliet Java and 70% on Juliet C/C++ respectively. GPT-4 and CL-34B perform extremely well on Out-of-bounds Read (78.6% and 64.3%) and Out-of-bounds Write (78.1% and 71.9%) on the CVEFixes Java dataset.
@@
Poor performance on real-world C/C++ is due to missing global context. We see that the performance of all LLMs on vulnerabilities in CVEFixes C/C++ is worse than that on the same CWEs in CVEFixes Java and Juliet C/C++. For instance, while GPT-4 and CL-34B perform extremely well on the Out-of-bounds Read / Write vulnerabilities in CVEFixes Java as discussed above, they report accuracies ¡ 53% for these CWEs on the CVEFixes C/C++ dataset. We attribute this disparity to the nature of these vulnerabilities in the two languages: Out-of-bounds Reads / Writes in CVEFixes C/C++ require reasoning about pointers and structs which requires more context about the structs and their members. In CVEFixes Java, on the other hand, these vulnerabilities arise primarily due to illegal array indexing. This issue does not emerge in Juliet C/C++ because all the information about the pointers is presented in the snippet. We present examples in Appendix A.8.
Some LLMs are better at detecting certain CWEs. GPT-4 reports high accuracies on CWE-78 (OS Command Injection) in 3/5 datasets (CVEFixes C/C++ with 70.8%, CVEFixes Java with 61.7%, and Juliet Java with 63.8%) but the other models do not. Among the CodeLlama models, CL-13B performs well on CWE-362 (Race Condition) in CVEFixes C/C++ with 63.1% and CVEFixes Java with 66.7%. CWE-89 (SQL Injection) is another vulnerability where only CL-7B reports high accuracies on 2/3 datasets (Juliet Java with 74% and CVEFixes Java with 64%).
3.4. RQ4: Impact of Adversarial Attacks
Experiment setup. We implement three semantics-preserving attacks, proposed by Gao et al. (Gao et al., 2023): dead-code injection, variable renaming, and branch insertion (illustrated in Table 11 in appendix). We only select test samples where the original prediction of LLMs (with the CWE prompt) was correct. We select 100 test samples per dataset for this experiment and measure the reduction in accuracy after the attack. For this experiment, we select two models, GPT-4 and CL-7B, and four datasets. Figure 4 presents our results.
Results. We find GPT-4 suffers mild degradation in accuracy (average 8.6%) across all datasets and attacks. For GPT-4, the degradation is more significant for real-world datasets (up to 23%) compared to synthetic ones (up to only 2%). In contrast, for CodeLlama-7B, the accuracy significantly reduces by almost 40-50% in many cases and by 39% on average, especially for C++.
Among the three attacks, branch insertion (27% avg. reduction) and variable renaming (23% avg. reduction) have the highest impact on LLM performance. On further analysis of incorrect predictions, we observe that under these attacks, GPT-4 either fails to infer any sources, sinks, or sanitizers (even if they exist) or the reasoning chain breaks (which worked previously). Interestingly, under these attacks, CodeLlama-7B fails to follow the instructions in the system prompt and does not perform proper analysis using the dataflow analysis steps, causing it to output an incorrect response.
3.5. RQ5: LLMs vs Static Analysis Tools

### v3

@@ II-B Metrics
F1 score: The F1 score is a harmonic mean of precision and recall. It is computed as: .
II-C Large Language Models
We choose the most popular state-of-the-art pre-trained Large Language Models (LLMs) for our evaluation. We choose three closed-source models (GPT-4, GPT-3.5 and Gemini-1.5-Flash) and thirteen open-source models from the Codellama-x, Llama-3.1-x, Mistral-Codestral-x, DeepSeekCoder-x, Qwen2.5-x and Qwen2.5-Coder-x series. We use the “Instruct” variants of the models wherever applicable since they are fine-tuned to follow user instructions and hence can better adapt to specific reasoning tasks. We access the GPT-x models and Gemini-1.5-Flash using the OpenAI and Google Gemini APIs respectively and use the Hugging Face APIs [38] to access the open-source models. Table V presents more details about the models.
| Model | Model Version | Size | Context Size |
| Qwen-2.5C-1.5B | qwen2.5-coder-1.5b | 1.5B | 128K |
| Qwen-2.5C-7B | qwen2.5-coder-7b | 7B | 128K |
| CodeLlama-7B | Codellama-7b-instruct | 7B | 16K |
| DSCoder-7B | deepseekcoder-7b | 7B | 4K |
| Llama-3.1-8B | llama-3.1-8b | 8B | 128K |
| CodeLlama-13B | CodeLlama-13B-Instruct | 13B | 16K |
| Qwen-2.5-14B | qwen2.5-14b | 14B | 128K |
| DSCoder-15B | deepseekcoder-v2-15b | 33B | 128K |
| Codestral-22B | mistral-codestral-22b | 22B | 32K |
| Qwen-2.5-32B | qwen2.5-32b | 32B | 128K |
| DSCoder-33B | deepseekcoder-33b | 33B | 16K |
| CodeLlama-34B | CodeLlama-34B-Instruct | 34B | 16K |
| Llama-3.1-70B | llama-3.1-70b | 70B | 128K |
| Gemini-1.5-Flash | gemini-1.5-flash | N/A | 1M |
| GPT-3.5 | gpt-3.5-turbo-0613 | N/A | 4K |
@@ II-C Large Language Models
II-D Prompting Strategies for LLMs
We explore various prompting strategies that can assist LLMs in predicting if a given code snippet is vulnerable. The LLMs discussed in this study support chat interactions with two major types of prompts: the system prompt can be used to set the context for the entire conversation while user prompts can be used to provide specific details throughout the chat session. We include a system prompt at the start of each input to describe the task and expected structure of the response. Since persona assignment has been shown to improve the performance of GPT-4 on specialized tasks [39], we add the line “You are a security researcher, expert in detecting security vulnerabilities” at the start of every system prompt to assign a persona of a Security Researcher to the model. The system prompt for all experiments ends with the statement “Provide response only in the following format:” followed by an expected structure of the response from the model. The system prompt is followed by a user prompt that varies across the various prompting strategies. In all our experiments, we incorporate the target code snippet into the user prompt without any changes.
We construct three different prompting strategies:
II-D1 Basic prompt
We design a very simple prompt (shown in Listing 4 in the Appendix) to test if the model can take a target code snippet as input and detect if it is vulnerable and determine the correct CWE as well. The prompt begins with the message “Is the following code snippet prone to any security vulnerability?” followed by the code snippet.
| Steps | OWASP | Juliet | Juliet | CVEFixes | CVEFixes | Total |
@@ II-D2 CWE specific prompt
The CWE specific prompt is presented in Listing 5 in the Appendix. This prompt is similar to the Basic prompt except that it asks the model to predict if the given code snippet is vulnerable to a specific target CWE. Hence, the user prompt starts with “Is the following code snippet prone to CWE?” followed by the code snippet. For instance, for CWE-476, the user prompt would start with “Is the following code snippet prone to CWE-476 (NULL Pointer Dereference)?” followed by the target code snippet.
II-D3 Dataflow analysis-based prompt
@@
Dataflow analysis is used by several static analysis tools to infer if there exists an unsanitized path from a source to a target node. Further, prior literature has shown step-by-step instructions can often elicit better reasoning from LLMs [40]. Motivated by these observations, we designed the CWE-DF prompt (shown in Listing 6 in Appendix) that prompts the model to simulate a source-sink-sanitizer-based dataflow analysis on the target code snippet before predicting if it is vulnerable. Naturally, this prompt is costlier since it generates more tokens than the other prompts. We provide the full prompts in Appendix B.
II-D4 Other prompting strategies
@@
We also tried other prompting strategies such as Few-shot prompting and Chain-of-thought prompting. In the few-shot prompting setup, we include two examples of the task (one with a vulnerability and one without) in the CWE specific prompt before providing the target code snippet. With Chain-of-thought prompting, we prompt the model to generate a reasoning chain before the final answer by adding a “Let’s think step-by-step” statement at the end of the CWE specific prompt. Our initial experiments with GPT-4 prompted using these techniques did not yield results better than the Dataflow analysis-based prompt on 100 random samples from two datasets. Hence, we do not include these prompts in this study. We refer readers to Appendix C for more details.
II-E Dataset Processing and Selection
@@
We preprocess each dataset before evaluation by removing or anonymizing information such as commits, benchmark IDs, or vulnerability names that may provide obvious hints about the vulnerability. We also skip benchmarks that are spread across multiple files, due to limitations of prompt size. We only consider samples corresponding to vulnerability types (CWEs) listed in MITRE’s Top 25 Most Dangerous Software Weaknesses [41]. We then filter code snippets with more than 2048 tokens due to prompt size limitations and randomly sample 500 vulnerable and 500 non-vulnerable samples per dataset. Table VI presents the details of our selection stages. We provide more details for each dataset in Appendix A.
II-F Experimental Setup
Experiments with closed-source models. We use the OpenAI public API’s ChatCompletions endpoint to perform the experiments with GPT-3.5 and GPT-4. We use Google’s Gemini API for the experiments with Gemini-1.5-Flash.
Experiments with open-source models. We run all experiments with the open-source LLMs using the HuggingFace API on a cluster with A100, A6000, and RTX 2080 GPUs.
@@ II-F Experimental Setup
In all our experiments, we set the sampling temperature to 0 for obtaining deterministic predictions, the maximum number of tokens to 1024, and use the default values for all other parameters. We use the top-1 predictions for evaluation.
III Results
III-A RQ1: Effectiveness of LLMs
@@
We evaluate the performance of pre-trained LLMs on five open-source datasets discussed in Section II-A. Figure 1 presents the best accuracy and F1 scores (across prompts) of the 16 models from Table V on all datasets. The more detailed metrics for all prompts are presented in Appendix D.
Finding 1.1: Modest Vulnerability Detection Performance Across LLMs. The best performing models and prompts per dataset report an accuracy of 62.8% on average. The highest of these is reported by Llama-3.1-70B (with CWE) on the Juliet Java dataset (76%). The other best performing models per dataset are: CodeLlama-13B on OWASP (60%), Gemini-1.5-Flash on CVEFixes Java (57%), Qwen-2.5-14B on Juliet C/C++ (65%) and Qwen-2.5-32B on CVEFixes C/C++ (56%). This confirms that no model individually performs the best across multiple datasets. Moreover, the best accuracies on synthetic datasets are 10.5% higher on average than those on the real-world datasets, suggesting that these models might not be suitable for real-world vulnerability detection yet.
Finding 1.2: Performance does not improve with scale. On the real-world datasets, Qwen-2.5-14B and Qwen-2.5-32B report higher accuracies than the GPT-x models despite being much smaller. We see many similar patterns across all models studied implying that model scale does not impact performance. Within the same model class, the GPT-x models and the Llama-3.1-x models exhibit improvements in accuracy as the size of the model increases on at least three datasets. However, this is not observed in other model classes, i.e., the Qwen-2.5C-x, Qwen-2.5-x, Codellama-x and the DSCoder-x series. This is in stark contrast to other domains where increasing model size leads to better performance. Listing LABEL:lst:juliet1724 presents a representative example where CodeLlama-7B correctly predicts that an integer overflow vulnerability (CWE-190) cannot occur in the given context while CodeLlama-13B does not.
@@
III-B RQ2: Comparison of Prompting Strategies
@@
Figure 2 presents the accuracy and F1 scores (averaged across all LLMs) of the three prompting strategies across all datasets. Overall, the three prompts perform similarly in terms of accuracy on all datasets with CWE-DF reporting slightly higher accuracies on the real-world datasets. Interestingly, the CWE-DF prompt reports significantly higher F1 scores on average than CWE and Basic prompt on the real-world datasets (by upto 0.18 on CVEFixes C/C++ and 0.14 on CVEFixes Java). Furthermore, CWE-DF reports the lowest variance in accuracies and F1 scores across models, as suggested by the standard deviation bars in Figure 2. On the other hand, the Basic prompt reports the highest variance in F1 scores. We next highlight qualitative differences between prompts:
@@
Finding 2.1: Specifying the CWE in the prompt reduces false alarms. Table VII presents the percentage of samples where the Basic prompt predicts the correct CWE, averaged across datasets for 4 models. We observe that the Basic prompt detects incorrect CWEs in 60% and 53% of all Java and C/C++ samples across models. We further manually inspected 10 vulnerable and 10 non-vulnerable samples from Juliet Java where GPT-4 with Basic prompt is incorrect. In only 4 or the 20 samples, the Basic prompt predicts a plausible CWE. However, these CWEs are unlikely due to the context. For example, it predicts that a value read from an input stream can be vulnerable if not validated (CWE-20) but this value is not used in a vulnerable context. The CWE specific prompt (i.e., the Basic prompt with CWE) improves or retains accuracy over the Basic prompt on all 5 datasets. GPT-4 with the CWE specific prompt not only correctly predicts all the 20 samples discussed above but also provides useful high-level explanations for why the snippet is vulnerable/not vulnerable in 18 / 20 samples. The 2 incorrect explanations are artifacts of faulty reasoning or hallucination: for example, an Integer Overflow due to addition to INT_MAX in the function is incorrectly attributed to subtracting from INT_MIN in the explanation. Based on these observations, specifying the CWE in the prompt can be helpful in reducing incorrect predictions.
| Language (Avg.) | GPT-4 | GPT-3.5 | CodeLlama-34B | CodeLlama-13B | |
| Java | 0.41 | 0.34 | 0.37 | 0.38 | |
| C/C++ | 0.29 | 0.31 | 0.33 | 0.35 |
@@
Finding 2.2: Dataflow analysis identifies CWE-relevant textual cues and provides actionable explanations. The Dataflow analysis-based prompt (CWE-DF) performs better than CWE specific prompt on the real-world datasets, CVEFixes C/C++ and CVEFixes Java. We inspect 10 vulnerable and 10 non-vulnerable samples from CVEFixes Java where GPT-3.5 correctly predicts only with CWE-DF. We find that the CWE-DF prompt leverages textual cues for sanitization (e.g., csrftokenhandler() suggests protection from CSRF) in 16/20 samples that are missed by CWE specific prompt.
Further, CWE-DF responses are more useful in localizing the vulnerability as it correctly predicts the correct sources and sinks in 18 / 20 samples, sanitizers in 16 / 20 samples, and unsanitized flows in all samples. Listing LABEL:fig:motivating-example presents a response from GPT-4 using CWE-DF prompt that correctly identifies the unsanitized flows between sources and sinks. We present more CWE-DF examples in Appendix F.
Finding 2.3: CWE-DF identifies the correct sources and sinks even when the final prediction is incorrect. We also inspect 10 vulnerable and 10 non-vulnerable samples from Juliet C/C++ where CWE-DF’s predictions are incorrect. Surprisingly, the sources, sinks and sanitizers are correctly identified in 17 / 20 samples but the unsanitized flows are incorrect in 12 / 20 samples. Hence, the final predictions are incorrect only due to erroneous reasoning about the snippet / false assumptions about the CWE. Listing LABEL:fig:julietcpp43583 presents an example where the vulnerability is not detected but the sources and sinks are correctly identified. This suggests that the CWE-DF prompt can be used to accurately identify sources/sinks/sanitizers and other dataflow analysis techniques can be used to reason about unsanitized flows to predict the vulnerability.
@@
III-C RQ3: Performance of LLMs across CWEs
@@
We next evaluate how the LLMs perform on different classes of security vulnerabilities (CWEs). Because the CWE-wise distribution of vulnerable and non-vulnerable samples can be imbalanced, we compute balanced accuracy for each CWE (for ease of presentation, we refer to it as accuracy henceforth in this section). For each dataset and model, we consider the best-performing prompt for the analysis and only report CWEs with at least 10 samples. Figure LABEL:fig:java_cwe_distribution presents the CWE-wise distribution of accuracies on the real-world datasets, CVEFixes Java and CVEFixes C/C++. Figure 3 reports the accuracies on the synthetic datasets, OWASP, Juliet Java and Juliet C/C++.
Finding 3.1: LLMs perform well on vulnerabilities that do not require additional context. These CWEs include OS Command Injection (CWE-78), Out-of-bounds Read / Write (CWE-125, CWE-787), Null Pointer Dereference (CWE-476), Cross-site Scripting (CWE-79), SQL Injection (CWE-89) and Incorrect Authorization (CWE-863). The higher performance on these vulnerabilities can be attributed to the fact that these are fairly self-contained and little additional context is needed to detect them. For example, 4/16 LLMs report accuracies 60% on Incorrect Authorization (CWE-863) on CVEFixes Java it is easier to validate if an implemented authorization check is correct or not. On the other hand, no LLMs report high accuracies on Missing Authorization (CWE-862) since it’s not known if an input parameter has been authorized outside the context of the target method and more context is hence needed to detect this vulnerability class. The following summarizes how LLMs perform on these CWEs:
OS Command Injection (CWE-78) sees the highest performance across models and datasets with 60% accuracies reported by 5/16 LLMs on CVEFixes Java and 11/16 LLMs on CVEFixes C/C++. Llama-3.1-70B reports the best performance on CWE-78 with accuracies 64% and 78% on CVEFixes Java and CVEFixes C/C++ respectively.
Llama-3.1-8B and GPT-4 perform extremely well on Out-of-bounds Write (CWE-787) with accuracies of 89% and 79% on CVEFixes Java and CVEFixes C/C++ respectively and on Out-of-bounds Read (CWE-125) with accuracies of 84% and 78% respectively. Moreover, accuracies over 60% are reported on CVEFixes Java by 5/16 LLMs for CWE-125 and 4/16 LLMs for CWE-787.
NULL Pointer Dereference (CWE-476) sees accuracies higher than 60% by 3/16 LLMs on CVEFixes C/C++, 7/16 on Juliet C/C++ and 11/16 on OWASP. Interestingly, DSCoder-7B performs the best on all three datasets with accuracies of 63%, 88% and 83% respectively.
GPT-3.5 reports the highest accuracy of 70% on SQL Injection (CWE-89) on CVEFixes Java and 7/16 LLMs report accuracies over 60%. Surprisingly, all LLMs report accuracies 60% on the same CWE on synthetic datasets (OWASP and Juliet Java).
Qwen-2.5-32B reports high accuracies of 67% and 66% on Cross-Site Scripting (CWE-79) on OWASP and CVEFixes Java respectively. Accuracies over 60% are reported by 3/16 LLMs on CVEFixes Java and 11/16 LLMs on OWASP.
@@
Finding 3.2: Poor performance on real-world C/C++ is due to missing global context. We see that the performance of all LLMs on vulnerabilities in CVEFixes C/C++ is worse than that on the same CWEs in CVEFixes Java and Juliet C/C++. While GPT-4 and Llama-3.1-8B perform extremely well on the Out-of-bounds Read / Write vulnerabilities in CVEFixes Java as discussed above, they report accuracies 53% for these CWEs on the CVEFixes C/C++ dataset. In fact, no LLMs report accuracies 60% for these CWEs on CVEFixes C/C++. We attribute this disparity to the nature of these vulnerabilities in the two languages: Out-of-bounds Reads / Writes in CVEFixes C/C++ require reasoning about pointers and structs which requires more context about the structs and their members. In CVEFixes Java, on the other hand, these vulnerabilities arise primarily due to illegal array indexing. This issue does not emerge in Juliet C/C++ because all the information about the pointers is presented in the snippet. We present examples in Appendix G.
Finding 3.3: Some LLMs are better at detecting certain CWEs. Concretely, the following LLMs report the best accuracies on certain CWEs across datasets:
Llama-3.1-70B on OS Command Injection (CWE-78)
DSCoder-7B on NULL Pointer Dereference (CWE-476)
Qwen-2.5-32B on Cross-Site Scripting (CWE-79)
Llama-3.1-8B and GPT-4 on Java Out-of-bounds Read/Write (CWE-125/787)
III-D RQ4: LLMs vs Static Analysis Tools